Sorry, you need to enable JavaScript to visit this website.

MCST Banner Image

am1weekly_2001-027_2001-033

Mission Operations Days: 2001/027 to 2001/033
January 27, 2001 00:00:00 GMT to February 02, 2001 00:00:00 GMT


Terra Spacecraft and MODIS Instrument Status:

Terra (AM-1) is in Normal Ops Mode
MODIS is in Science Mode - B Side
ComponentState(s)Comment
BlackbodyA Off; B On (290 K)Nominal
Calibration ElectronicsA Off; B OnNominal
Control ProcessorA Off; B OnNominal
Nadir DoorUnlatched, OpenNominal
Space View DoorUnlatched, OpenNominal
Solar Diffuser DoorUnlatched, ClosedNominal
FDDI FormatterA Off; B OnNominal
FDDI PortA On; B OffNominal
FIFO Memory#1 & #2 Off; #3 & #4 OnNominal
Format ProcessorA Off; B OnNominal
Power Supplies#1 Off; #2 OnNominal
PV VIS FPAA Off; B OnNominal
PV NIR FPAA Off; B OnNominal
PV SMIR FPAA Off; B OnNominal
PV LWIR FPAA Off; B OnNominal
PC LWIR FPAA Off; B OnNominal
RC Outgas HeatersAll Off (CS, IS, OS)Nominal
RC LWIR FPA HeaterOn (83 K)Nominal
RC SMIR FPA HeaterOffNominal
Scan AssemblyA Off; B OnNominal
SDSMA Off; B OffNominal
SRCAA Off; B OffNominal
PS1 Survival HeaterEnabledNominal
PS2 Survival HeaterEnabledNominal
Timing GeneratorA Off; B OnNominal
Flight SoftwareRev BD + 1 set of patches
  • FR Error Counting and Statistics (2 patches)
Nominal
Inhibit Ids SetNoneNominal
TMONs Enabled68Nominal

This Week’s Completed MODIS Activities:

Saturday, January 27, 2001

None

Sunday, January 28, 2001

None

Monday, January 29, 2001

None

Tuesday, January 30, 2001

None

Wednesday, January 31, 2001

2001/031 20:49:46 OA-16: SDSM SCREENED
2001/031 22:28:39 OA-15: SDSM OPEN

Thursday, February 01, 2001

None

Friday, February 02, 2001

None

This Week's Scheduled MODIS Activities Not Completed:

None

Upcoming MODIS Events:

Friday, February 02, 2001

None

Saturday, February 03, 2001

None

Sunday, February 04, 2001

None

Monday, February 05, 2001

None

Tuesday, February 06, 2001

None

Wednesday, February 07, 2001

None

Thursday, February 08, 2001

2001/039 15:50:51 – 16:24:24 ATC OA-19: SRCA Full Radiometric

Friday, February 09, 2001

2001/040 14:52:39 – 15:44:18 ATC OA-22: SRCA Full Spectral (Part 1)
2001/040 16:48:22 – 17:23:07 ATC OA-22: SRCA Full Spectral (Part 2)
2001/040 18:10:24 – 19:01:40 ATC OA-22: SRCA Full Spectral (Part 3)
2001/040 20:04:25 – 20:42:23 ATC OA-22: SRCA Full Spectral (Part 4)

Maneuvers:

A drag makeup maneuver was performed by the FOT at 017/19:38:00 to 017/21:08:00.

The next MODIS Lunar Calibration Roll Maneuver is planned for 2001/043, February 12th.

MODIS Anomalies:

No formatter anomalies have been detected since switching to B-side science operations. The MODIS IOT is continuing to monitor changes with the radiative cooler temperatures and focal plane power.

General Instrument Comments:

MODIS is in Science Mode on the B-side with the SVD and NAD open, operating with the Formatter Resolution Flight Software Patches (FRFSP). No new events have been witnessed by the FRFSP since the B-Side Transition on DOY 2000/304.

MODIS Telemetry Trends:

The Radiative Cooler intermediate stage temperature has slowly been increasing since approximately day 2000/240. During this time the LWIR focal plane heater margin has been decreasing. The IOT is continuing to trend these telemetry mnemonics.

Temperature trends are correlating with the solar beta angle. This includes a steadying of the intermediate stage outgas temperature. This temperature is expected to increase again when the beta angle gets smaller.

Non-MODIS Significant Events:

None.

Limited Life Item Status:

SRCA 10W Lamp #1: 179.2 of 500 hours
SRCA 10W Lamp #2: 139.2 of 500 hours
SRCA 10W Lamp #3: 148.4 of 500 hours
SRCA 10W Lamp #4: 61.5 of 500 hours

SRCA 1W Lamp #1: 556.6 of 4000 hours
SRCA 1W Lamp #2: 276.3 of 4000 hours

Solar Diffuser Door: 1519 of 3022 Movements
Nadir Aperture Door: 532 of 1316 Movements
Space View Door: 437 of 1316 Movements